### v4

# MeSH Concept Relevance and Knowledge Evolution: A Data-driven Perspective

Abstract
The Medical Subject Headings (MeSH), one of the main knowledge organization systems in the biomedical domain, constantly evolves following the latest scientific discoveries in health and life sciences. Previous research focused on quantifying information in MeSH using its hierarchical structure. In this work, we propose a data-driven approach based on information theory and network analyses to quantify the relevance of MeSH concepts. Our approach leverages article annotations and their citation networks to compute informativeness, usefulness, disruptiveness, and influence of MeSH concepts over time. Using the the citation network and the MeSH hierarchy, different relevance aspects are computed, and an aggregation algorithm is used to propagate the relevance scores to the parent nodes. We evaluated our approach using changes in the terminology and showed that it effectively captures the evolution of MeSH concepts. At the first level of the hierarchy, the most relevant concept - Chemical and Drugs - had a decreasing trend (p-value $<0.01$), while at the second level, the most relevant concept - Neoplasms - had an increasing trend (p-value $<0.01$). We show that the mean relevance of evolving concepts is higher for concepts that remained unchanged (2.09E-03 vs. 8.46E-04). Moreover, we validated the ability of our framework to characterize retracted articles and showed that concepts used to annotate retracted articles (mean relevance: 0.17) differ substantially from those used to annotate non-retracted ones (mean relevance: 0.15). The proposed framework provides an effective method to rank concept relevance and can be useful in maintaining evolving knowledge organization systems.
keywords:
terminology evolution , concept relevance , information theory , network analysis
journal: Journal of Knowledge-Based Systems
[…]
[inst1]organization=University of Geneva,city=Geneva, postcode=1211, country=Switzerland
[…]
[inst2]organization=Université Paris-Saclay, CNRS, Laboratoire Interdisciplinaire des Sciences du Numérique,city=Orsay, postcode=91400, country=France
1 Introduction
The vast amount and complexity of the scientific biomedical literature require sophisticated methods of knowledge structuring and reasoning [1, 2]. Knowledge organization systems (KOSs), such as terminologies, thesauri, and ontologies, have been designed and used to encode concepts to improve discoverability, knowledge representation, and reasoning in biomedical corpora [1]. Among those KOSs, the Medical Subject Headings (MeSH) vocabulary indexes, structures, and enables search in the most prominent scientific biomedical corpus, i.e., the MEDLINE database accessible via the PubMed interface [3].
MeSH is a controlled vocabulary maintained by the National Library of Medicine [4]. In essence, it is a thesaurus that includes different types of concepts [1, 4, 5]. Descriptors are the primary elements of indexing and retrieval and are used to annotate the MEDLINE/PubMed dataset. Qualifiers and supplementary records are also available to provide additional information concerning descriptors. MeSH headings denote concepts from the biomedical domain and serve as a standardized interpretation for a group of medical terms [1, 4]. These headings are hierarchically organized into 16 main categories and can be associated with corresponding descriptors, which allows a direct mapping of descriptors to their appropriate MeSH headings. In addition to supporting search and question-answering in the biomedical literature [6, 7], they help explore patterns characterizing the structure and evolution of the biomedical field, assist in evaluating research grants and awards, and identify research topics of high interest [8, 9, 10].
As the knowledge in the biomedical domain evolves, MeSH needs to be periodically modified and updated to reflect these changes [11, 12]. Based on the current knowledge state, MeSH concepts (i.e., headings) encode a certain amount of information, which might vary over time [13]. The relevance of some concepts may decrease or increase, making them obsolete or requiring them to change drastically or to be added [5, 13, 14]. As a consequence, the hierarchy structure can change. Thus, it is crucial to identify the most representative MeSH concepts [1, 15] to summarize the content of the thesaurus, prioritize actions for maintaining tasks in KOS concepts, and enhance information retrieval systems relying on KOS, such as PubMed [14, 16, 17].
Different approaches have been used to quantify the amount of information encoded in MeSH by analyzing various features, such as the hierarchy, bibliometric indices of articles, and the co-occurrence of descriptors [9, 5, 13, 18, 19]. The hierarchy was used to examine properties, and MEDLINE articles citing descriptors were analyzed using scientometric methods [13, 18]. Balogh et al. discovered that nodes in this thesaurus with many children are more likely to gain even more children [13]. Adding new concepts to the hierarchy creates a non-uniform distribution among the primary branches, and adding links follows a non-trivial preference. Chien et al. analyzed PubMed articles in medicine and health using the x-index [18]. Descriptors were used to analyze trends in general and personalized medicine [9]. The co-occurrence of descriptors and new descriptors was also investigated [5, 19]. Kastrin and Hristovski built a network using co-occurrences of descriptors to investigate indexing trends [19]. Nentidis et al. demonstrated that about a quarter of new descriptors in this thesaurus can be considered emerging concepts [5]. In contrast, the remainder of the concepts were derived from descriptors in previous releases. In particular, Ilgisonis et al. focused on related descriptors rather than purely annotated descriptors, i.e., descriptors that co-occurred with the annotated descriptors [9]. Despite some limited attempts to use instances, such as the occurrence of descriptors in PubMed articles, instance analysis remains largely unexplored for the quantification of MeSH concepts over time.
To bridge this gap, this study aims to investigate an approach to measure the relevance encoded by MeSH considering the instantiation of its concepts. To do so, we analyze three main factors - the usage of concepts in PubMed articles, the hierarchical structure of MeSH, and the citation network of PubMed articles - in view of information theory and network analyses. The approach includes four aspects to quantify the relevance of MeSH concepts: informativeness, usefulness, disruptiveness, and influence. To quantify the amount of information a concept conveys, i.e., its informativeness, we use entropy, the standard concept of information theory [20, 21]. We use category utility supported by the MeSH hierarchical structure to measure the usefulness of a concept [22]. Some concepts may be more disruptive due to different factors, defining their disruptiveness level. For instance, they can be used to annotate highly cited but consolidated technologies, not bringing novelty to the field [10, 23, 24]. Lastly, to quantify the influence of MeSH concepts, we use the centrality metric from graph theory [25], which can be used to quantify node importance in networks [26]. We applied these methods to the MEDLINE database, analyzing their MeSH annotations over time. We generated monthly data by mapping references to MeSH concepts and computing each metric for each concept. This data was then propagated to the MeSH hierarchy. Finally, results were combined using a ranking fusion method - reciprocal ranking fusion (RRF) - to aggregate the multidimensional scores and provide a relative relevance of the MeSH concepts.
The paper presents three main contributions. First, we propose an algorithm to compute the relevance of a MeSH concept by propagating information from a network of data instances to the KOS structure. The network of MeSH instances is based on the PubMed citation network, while the knowledge structure relies on the thesaurus hierarchy. Second, we assess the relevance of MeSH concepts quantified by multidimensional metrics - informativeness, usefulness, disruptiveness, and influence - and show how they vary at different hierarchical levels over time. Lastly, we provide a quantitative evaluation of the proposed approach by analyzing changes in the MeSH terminology and retracted papers.
2 Materials and Methods
We explore the evolution of MeSH by analyzing concept instantiations over different years in PubMed. For this purpose, we use PubMed/MEDLINE corpus, which contains articles manually annotated with MeSH descriptors. In this section, we explain the methodology used to create the dataset[dataset] [27]. We describe the metrics used to quantify relevance in a KOS. Then, we discuss strategies for propagating metrics to the hierarchy and computation of a metrics fusion. To conclude this section, we detail the evaluation criteria used in this work.
2.1 Dataset and pre-processing
We conducted our study using the 2022 collection of PubMed [^1] https://lhncbc.nlm.nih.gov/ii/information/MBR/Baselines/2022.html, which includes 33,403,054 articles collected until December 2021 (see the distribution of PubMed articles in Figure 7). PubMed indexes new biomedical and life science articles whenever they are published. Domain experts annotate articles in PubMed using MeSH descriptors, which are organized hierarchically into 16 categories represented by single letters, such as A for Anatomy, B for Organisms, and C for Diseases. The MeSH hierarchy’s complete list of categories and their respective numbers of instances per MeSH version are presented in Table 1.
Table 1: Number of articles (in millions) that map to different categories in the MeSH hierarchy over time.
Code Category label 2014 2015 2016 2017 2018 2019 2020 2021
A Anatomy 9.6 10.0 10.4 10.7 11.1 11.5 11.9 12.3
B Organisms 19.7 20.7 21.6 22.5 23.5 24.4 25.4 26.3
C Diseases 11.6 12.2 12.7 13.3 13.9 14.5 15.1 15.7
D Chemicals and Drugs 12.8 13.3 13.9 14.4 15.0 15.5 16.1 16.6
E
Analytical, Diagnostic and
Therapeutic Techniques, and
Equipment
15.4 16.1 16.9 17.6 18.3 19.1 19.8 20.5
F Psychiatry and Psychology 3.5 3.8 4.0 4.2 4.4 4.6 4.8 5.0
G Phenomena and Processes 12.9 13.5 14.1 14.7 15.3 15.9 16.6 17.1
H Disciplines and Occupations 2.8 2.9 3.0 3.2 3.3 3.4 3.6 3.7
I
Anthropology, Education,
Sociology, and Social
Phenomena
2.2 2.3 2.5 2.6 2.7 2.9 3.0 3.2
J
Technology, Industry, and
Agriculture
1.9 2.0 2.2 2.3 2.4 2.6 2.7 2.8
K Humanities 0.5 0.6 0.6 0.6 0.6 0.7 0.7 0.7
L Information Science 2.5 2.7 2.8 3.0 3.1 3.2 3.4 3.5
M Named Groups 7.7 8.1 8.5 8.9 9.4 9.8 10.2 10.6
N Health Care 8.3 8.8 9.4 9.9 10.4 11.0 11.6 12.1
Z Geographicals 2.7 2.9 3.1 3.3 3.5 3.7 3.9 4.1
For this collection, we built citation networks using PubMed article citations and the [dataset]iCite [^2] https://nih.figshare.com/articles/dataset/iCite_Database_Snapshot_2022-04/19763386 dataset [28] for a more comprehensive source of citations. We then created accumulative monthly networks, i.e., including nodes from previous months, from January 2014 until December 2021. To build the corresponding MeSH instance hierarchy, we used the MeSH descriptors assigned to articles in a particular month mapped to their respective MeSH tree code (See Figure 1d for an example). The first monthly citation network (January 2014) has 18,728,651 nodes and 381,296,042 edges, and as citation networks accumulate articles over time, the last one (December 2021) has 27,728,971 nodes and 673,415,629 edges. Due to the prohibitive time required to compute certain metrics, we used 10% of nodes randomly sampled from the original monthly citation network, and edges were included only when both nodes existed in the sampled network. This resulted in the first sampled citation network (January 2014) having 1,872,866 nodes and 3,807,856 edges, while the last sampled network (December 2021) had 2,772,898 nodes and 6,745,896 edges.
2.2 MeSH relevance quantification
Analogous to the relevance definition in information science and information retrieval [29], we define the relevance of a concept as how well a concept or set of concepts meets the representation needs of a dataset [30]. Similar to the definition of relevance provided by Saracevic [31], in our study, we consider multiple aspects to quantify the relevance of MeSH concepts, namely informativeness, usefulness, disruptiveness, and influence (Figure 1a) [^3] https://zenodo.org/records/13635979. Each article in MeSH is annotated with MeSH descriptors. Furthermore, descriptors map to one or more tree nodes (Figure 1c and Figure 1d). In the following, we describe these metrics and how the relevance of a MeSH concept is quantified according to them.
Figure 1: Aspects to quantify the relevance of MeSH concepts. a) Metrics quantifying each aspect. b) Intuition of computation of each metric in this work. c) Flow of mappings over time: PubMed articles are annotated with MeSH descriptors, which link to one or more nodes in the MeSH hierarchy. d) Example of mapping between MeSH descriptors and the MeSH hierarchy of the article “Approaches towards fighting the COVID-19 pandemic (Review).” Bottom: MeSH descriptors. Top: MeSH hierarchy of the example’s article, showing only relevant nodes for this case.
2.2.1 Informativeness
To quantify the relevance of a concept in terms of informativeness, we use the entropy measure introduced by Shannon [20, 21]. Entropy quantifies the level of uncertainty of a variable. In our case, if a concept is used frequently to annotate multiple papers, it will not be informative, as many other articles also contain this annotation. For example, the concept “Humans” (D006801) is used to annotate approximately 70% of PubMed articles. Therefore, this concept adds little information when it appears in an article, having a low entropy value. On the other hand, the concept “Proteins” (D011506) is used to annotate only about 2% of PubMed articles, which makes it much more informative.
As shown in Equation 1, we follow Shannon’s entropy definition $H(X)$, where $p(x)$ represents the probability of an event’s occurrence and $x$ refers to all possible values that a variable $X$ can take:
$H(X)=-\sum_{x}p(x)\log_{2}p(x).$ (1)
In our study, an event is represented by the instantiation of a MeSH concept, that is, by articles being mapped to nodes in the MeSH hierarchy, representing the concept’s usage (dotted red arrows in Figure 1). The probability of an event $p(x)$ is then calculated as the probability that a concept $x$ is mapped within all mappings in its hierarchical level. Mappings in the lower levels are propagated to their parents when calculating the latter’s probability.
2.2.2 Usefulness
To quantify relevance in terms of usefulness, we use the category utility measure to identify basic-level categories in a taxonomy [22]. Basic-level categories are preferred categories in natural and artificial concept hierarchies, and they carry most information about the category’s attributes [22, 32] (e.g., when shown an image of a particular oncological exam, most professionals will identify it as a cancer rather than as a disease or a colon cancer). A node in a hierarchy is useful to the extent that it can be expected to improve the ability of a person to accurately predict the features of instances of that category. As shown in Figure 1b, in this study, the mapping of articles to the hierarchy characterizes concepts. Therefore, a concept that belongs to a highly mapped branch in the hierarchy will be more useful (e.g., as node c15 in the example).
Category utility combines the frequency of a category’s occurrence with the probability of its features [22]. Let us define a category $c$ as a finite set of instances: $c=\{o_{1},o_{2},...,o_{n}\}$, where each instance $o_{i}$ can be described by a finite set of discrete features $F=\{f_{1},f_{2},...,f_{m}\}$. The category utility (CU) of category $c$ with a set of features $F$ can then be calculated as in Equation 2, where $p(c)$ is the probability of a category (i.e., of a MeSH concept in our case), $p(f_{k})$ is the probability of feature $f_{k}$ and $p(f_{k}|c)$ is the conditional probability of feature $f_{k}$ given category $c$:
$CU(c,F)=p(c)\sum_{k=1}^{m}[p(f_{k}|c)^{2}-p(f_{k})^{2}].$ (2)
To calculate category utility for the nodes in the MeSH hierarchy, we use mapping to articles $a_{i}$ as the set of features $F=\{a_{1},a_{2},...,a_{m}\}$, where each feature $a_{i}$ has a binary value (annotated or not with a category). To compute the probabilities, we then created a binary matrix $M_{ij}$ of $n$ tree nodes and $m$ features as follows:
$M_{ij}=\begin{cases}1,&\text{tree node $i$ is mapped by article $j$}\\ 0,&\text{otherwise}\end{cases}.$
2.2.3 Disruptiveness
To measure the relevance of a concept in terms of the disruptiveness aspect, we use the disruption index [23]. This index was originally designed to quantify the extent to which an invention (a concept in our case) consolidates or destabilizes the subsequent use of the components (referenced concepts in our case) on which it builds [23]. To compute the disruption index $D$ for a focal node, we can consider the citation network as a tripartite graph $G=(V_{1},V_{2},V_{3},E)$, where the set of nodes ${V_{1},V_{2},V_{3}}$ represents the articles, which are connected by directed edges $E$, representing their citations [25, 26]. Given a focal node $a$ (light red in Figure 1b), corresponding to an article for which we want to compute the level of disruptiveness, there are $j$ articles that reference article $a$ and at least one of its predecessors (purple box), $i$ articles reference $a$ but none of its predecessors (green box), and $k$ articles do not reference $a$ but at least one of its predecessors (dark gray box), disruption $D$ for $a$ can then be defined by Equation 3:
$D=\frac{n_{i}-n_{j}}{n_{i}+n_{j}+n_{k}},$ (3)
where $n_{i}$, $n_{j}$, and $n_{k}$ represent the number of $i$, $j$, and $k$ articles, respectively. The disruption score $D$ of an article is then propagated to the concepts mapped to it ($c2$, $c3$, and $c4$ in the example). Finally, the disruption index $D_{c}$ for a concept $c$ is the sum over the disruption scores for all articles $a_{1}$, , $a_{n}$ which are mapped to $c$ divided by the number of articles in the citation network, i.e., $D_{c}=\sum_{i=1}^{n}D_{i}/m$, where $D_{i}$ is the disruption index of the $i$th article mapped to $c$ and $m$ is the number of articles in the citation network. To compute the $D$ index, we used Figueiredo and Andrade’s disruption implementation [23].
2.2.4 Influence
In graph theory, the influence of a node is quantified through graph centrality, which is based on the importance of its neighboring nodes. We use this concept to measure the influence aspect of a MeSH concept. Instead of considering the MeSH hierarchy, which captures only ontological relations, we use the instances (i.e., mapped articles) to identify the influence of MeSH concepts. Similar to the disruptiveness aspect, we consider the citation network to compute centrality. In Figure 1b, we show a small citation network example to demonstrate how nodes (articles) influence their neighbors. An article, annotated with one or more concepts, denoted by $c_{1}$, ..., $c_{n}$, gains more influence based on the influence of its incoming articles (e.g., light red box).
In our work, we utilized the PageRank algorithm to compute the centrality of a node [33, 25], which uses the importance of the node’s neighbors to compute the node’s importance. The mathematical definition of PageRank of node $i$ in a directed graph $G$ is $x_{i}$ provided by Equation 4:
$x_{i}=\alpha\sum_{j}A_{ij}\frac{x_{j}}{k_{j}^{out}}+\beta,$ (4)
where the parameter $\alpha$ regulates the contribution between the eigenvector and the constant term $\beta$. The term $\beta$ is used to initialize nodes with zero out-degree. The adjacency matrix of $G$ is denoted by $A_{ij}$, which captures the edges between pairs of nodes (citations in our case). The out-degree of node $j$ is represented by $k_{j}^{out}$. Similarly to disruptiveness, the centrality $x_{i}$ of an article $i$ is propagated to the concepts it is mapped to, and the overall centrality $x_{c}$ of a code is computed by summing over all the centralities of articles that concept $c$ is mapped to divided by the number of articles in the citation network, i.e., $x_{c}=\sum_{i=1}^{n}x_{i}/m$, where $x_{i}$ is the centrality score of the $i$th article mapped to $c$ and $m$ is the number of articles in the citation network. We used the implementation PageRank algorithm from the NetworKit [^4] https://networkit.github.io/ library to calculate the graph centrality, where $\alpha=0.85$ and $\beta=1-\alpha$.
2.3 Propagation of data instances to the knowledge representation
For the different concept relevance aspects, we encoded the usage of MeSH instances and propagated them to the knowledge representation. Given the “is a” relationship between MeSH concepts, we have considered two methods for propagating information to higher-level nodes: direct mapping and a cumulative weighted propagation algorithm. For the information-theoretical measures (entropy and category utility), which are computationally efficient, we propagate the mapping of instances directly to their parent concepts and compute the relevance scores for each node in the hierarchy. For example, if a paper is annotated with the MeSH concept “Heart Diseases [C14.280]”, its parents “Cardiovascular Diseases [C14]” and “Diseases [C]” are also considered to be mapped to that paper. Then, the information-theoretical metrics are computed normally for the parent nodes (e.g., C14 and C), including their direct mappings as well as the propagated mappings based on the “is a” relationship.
We approach the network analysis metrics (disruption and graph centrality) differently due to their high computational complexity. Instead of combining the mappings into higher-level nodes, we aggregate scores of lower-level nodes into the scores of their parents. The assumption is that if a node is disruptive or influential, so will its parent. As described in Algorithm 1, the propagation method requires the MeSH hierarchy $H$ and the network-based relevance scores computed using the citation network (variable $tnm$) as input. In line 2, we obtain tree nodes per level of specificity using the hierarchy $H$; for instance, tree nodes C and D are in level 1, while tree nodes C01 and D01 are in level 2. Then, we iterate over the hierarchy from bottom to top, using a reversed level list (variable $llr$, line 3) to propagate the relevance scores of the lower-level concepts to the higher levels. First, at each level, we get a set of unique parents (lines 5-9). Then, we retrieve their children (line 11) and, utilizing their initial measurements (lines 14-18), compute the weighted parent relevance, variable $gmp$, based on the children’s relevance and the number of nodes in their level (lines 19 and 21). Finally, the parent value is calculated taking into account the direct mapping to its node in addition to the weighted children measurements (lines 22 and 23).
Algorithm 1 Propagation of graph-based metrics to the MeSH hierarchy
0:  $H$ (hierarchy), $tnm$ (graph metric values initialized from descriptors)
0:  $gmh$ (graph metric of tree nodes in the hierarchy)
1:  $gmh\leftarrow\{\}$
2:  $dpt\leftarrow get\_nodes\_in\_levels\_specificity(H)$
3:  $llr\leftarrow\text{sorted list of keys in }dpt\text{ in reverse order}$
4:  for each $current\_level$ in $llr$ do
5:     $unique\_parents\leftarrow\text{an empty set}$
6:     for each $treenode$ in $dpt[current\_level]$ do
7:        $parent\leftarrow\text{get\_parent}(H,treenode)$
8:        add $parent$ to $unique\_parents$
9:     end for
10:     for each $parent$ in $unique\_parents$ do
11:        $children\leftarrow\text{get\_children($H$, $parent$)}$
12:        $gmp\leftarrow 0$
13:        for each $child$ in $children$ do
14:           if $child\in tnm$ then
15:              if $child$ has no children then
16:                 $gmh[child]\leftarrow tnm[child]$
17:              end if
18:           end if
19:           $gmp\leftarrow gmp+gmh[child]$
20:        end for
21:        $gmp\leftarrow gmp/\text{l}en(dpt[current\_level]$)
22:        $dgm\leftarrow tnm[parent]\text{ if }parent\in tnm\text{ else }0$
23:        $gmh[parent]\leftarrow gmp+dgm$
24:     end for
25:  end for
26:  return  $gmh$
2.4 Consolidating the relevance ranking of MeSH codes
Each of the metrics described previously measures different relevance aspects of a MeSH concept according to its data instances. To have a unified relevance score, we propose to fuse the multidimensional aspects into a single rank. Inspired by ranking fusion methods in the information retrieval field, we consider that each concept $d\in D$, where $D$ is the set of MeSH codes, is ranked according to a relevance dimension $r\in R$, where $R$ is the set of four aspects considered in this work: informativeness, usefulness, disruptiveness, and influence.
The field of information retrieval proposes various techniques for merging rankings, such as CombSUM, CombMNZ, CombANZ, rank-biased precision, and reciprocal rank fusion (RRF) [34]. Due to its simplicity, effectiveness, and robustness  [34, 35], we use RRF for computing the overall relevance of a MeSH concept based on their instantiations. RRF is an unsupervised approach that combines the outcomes of different information retrieval methods with varying relevance scores into a single result set using the rank information. The mathematical definition of RRF is provided in Equation 5:
$RRF\_score(d\in D)=\sum_{r\in R}\frac{1}{k+r(d)},$ (5)
where, in our case, a concept $d$ belongs to the set of MeSH codes $D$, and $r(d)$ represents the rank of a code $d$ according to a metric $r$ from the set of metrics $R$. The constant $k$ minimizes the impact of outliers.
2.5 Evaluation criteria
We qualitatively analyse the trends of MeSH concepts according to the different relevance aspects for various levels of the hierarchy and versions from 2014 to 2021. To quantitatively evaluate the effectiveness of our approach, we conducted two experiments. First, we assess the relevance of concepts that changed compared to those that remained unchanged between MeSH versions. Second, we compared the relevance of concepts between retracted and non-retracted papers in PubMed. We used the Mann-Whitney test to compare relevance changes; p-values $<$ $0.05$ are considered statistically significant.
3 Results
In this section, we provide the results of assessing the relevance of MeSH concepts according to the multi-aspect model previously described. We start with the analysis results of the MeSH terminology evolution, highlighting the most relevant concepts at the different levels of the hierarchy. Then, we provide quantitative analyses of concept relevance in two use cases: i) terminology evolution and ii) paper retraction. For the first, we compare the relevance of concepts that changed from one version to another versus those that remained unchanged. For the latter we compare the relevance of concepts used to annotate retracted papers versus those used to annotate non-retracted papers.
3.1 MeSH concept relevance trends
We calculated the relevance of MeSH concepts from 2014 to 2021 using the PubMed 2022 snapshot annotated with the MeSH 2022 version. Relevance scores for the different aspects - disruptiveness, influence, informativeness and usefulness - as well as for the overall concept relevance were computed monthly. Figure 2a depicts the cumulative trends (i.e., aggregate) for the first level of the MeSH hierarchy (A, …, Z). Disruptiveness, influence, and usefulness decrease over time ($p-value<0.001$). Park and colleagues also corroborate the decreasing trend of disruptiveness, although their analysis was based on scientific articles, in contrast to ours, which is based on MeSH concepts [36]. On the other hand, informativeness, which measures the degree of uncertainty, tends to increase over time ($p-value<0.001$). Therefore, the degree of uncertainty also increases as more recent knowledge is added to PubMed.
Figure 2: Trend of metrics relevance of MeSH concepts at level 1 and correlation of descriptors. a) The sum of MeSH categories per relevance aspect over time. b) MeSH categories per relevance aspect over time. c) Ranking of concepts at level 1 over time. d) Correlation of aspects. Column names are dis: disruptiveness (disruption), infl: influence (centrality), info: informativeness (entropy), use: usefulness (category utility), rel: concept relevance (RRF fusion). e) Trend of top concepts in categories C, D, and E at level 2 per metric.
Figure 2b shows how the first level of concepts in MeSH varies over time for each relevance aspect. All of the categories show a decreasing trend ($p-value<0.001$). The most useful concepts (Figure 2b - bottom right) in the first level are D (Chemicals and Drugs), followed by N (Health Care) and C (Diseases), indicating they encode the basic level categories of PubMed. In influence, all categories show a decreasing trend ($p-value<0.001$). Concept C is the most influential (Figure 2b - top right), indicating that ‘diseases’ is a central concept in PubMed. It is followed by the concepts D and G (Phenomena and Processes), which include biological, chemical, and physiological phenomena descriptors.
There is a group of non-medical concepts at the bottom in the trend of influence, grouping categories I (Anthropology, Education, Sociology, and Social Phenomena), N, H (Disciplines and Occupations), K (Humanities), M (Named Groups), L (Information Science), J (Technology, Industry, and Agriculture), Z (Geographicals) which present a more linear trend near 0. Trends of categories in disruption are mainly decreasing ($p-value<0.001$), though categories M and H are not statistically significant. J shows an increasing trend ($p-value<0.001$). On the other hand, K presents a growing trend, and I present no trend with no statistically significant $p-value$. Concept C is also the most disruptive in the first level of the MeSH hierarchy (Figure 2b - top left), being more than two times more disruptive than the top-2 concept (D). On the other hand, non-medical categories I, N, H, K, M, L, J, and Z are located at the bottom of this figure in an almost linear trend near 0, similar to the influence aspect. Trends of entropy are primarily increasing with a $p-value<0.001$. However, categories H, G, B, A, and D present a decreasing trend ($p-value<0.001$). Category C shows no trend (not statistically significant). Concept D has the highest informativeness score (Figure 2b - bottom left) followed by concept E (Analytical, Diagnostic and Therapeutic Techniques, and Equipment). However, it gradually decreases, with a standard deviation of 0.003377. The trends of this aspect remain constant over time, but this is due to the range of values of each MeSH category.
Figure 2c shows the ranking of concepts in level 1. Most of the concepts retain their ranking over time, except for Diseases (C), Health Care (N), Disciplines and Occupations (H), and I, which show some variation. Chemicals and Drugs (D) consistently ranks higher than other MeSH categories. The second position in this ranking alternates between ‘Diseases’ and ‘Health Care’, with the latter prevailing over the former for five out of eight years.
Figure 2d shows the correlation between relevance aspects for each descriptor using the monthly values in releases. The overall concept relevance, as a fusion of disruptiveness, influence, informativeness, and usefulness, shows a moderate correlation. Usefulness has a low correlation with disruptiveness and informativeness and a moderate correlation with influence. Influence and informativeness exhibit a moderate correlation.
Figure 2e shows the top concept trend in C, D, and E categories per metric, i.e., at level 2 in the hierarchy. Most of the trends are decreasing over time. For disruptiveness, concepts C04 (Neoplasms), D10 (Lipids), and E04 (Surgical Procedures) decrease over time ($p-value<0.001$). In influence, concepts C04, D05 (Macromolecular Substances), and E04 have a decreasing trend with a $p-value<0.001$. Concerning informativeness, C23 (Pathological Conditions, Signs and Symptoms) and E05 (Investigative Techniques) show an increasing trend, while D12 (Amino Acids, Peptides, and Proteins) has a decreasing trend ($p-value<0.001$). In usefulness, concepts C12 (Urogenital Diseases), D12, and E06 (Dentistry) present decreasing trends with a $p-value<0.001$. The increasing trend of concept C23 may mean that the manifestation of diseases is changing over time, producing uncertainty. On the other hand, the decreasing informativeness trend of concept D12 can mean that knowledge in this tree of MeSH is more established as this concept is used more frequently, becoming less informative. The trend of concept E05 shows an increasing degree of uncertainty until 2019 when the uncertainty decreases. The full description of the concepts, i.e., MeSH headings, can be found in Table 3
3.2 Top concepts in the MeSH terminology
Figure 3a displays the top 10 concepts according to the concept relevance, i.e., RRF, at level 2 (left side), level 3 (center), and level 4 (right). Half of the top concepts maintain their ranking over time. Concept G02 (Chemical Phenomena) shows variation among several positions, starting at the fourth position in 2014 and ending at the eighth in 2021. Meanwhile, C12 (Urogenital Diseases), D01 (Inorganic Chemicals), D02 (Organic Chemicals), and C01 (Infections) only change one position over time.
Figure 3: Top/bottom concepts over time per level 2 (left), 3 (center), and 4 (right). The left side of each figure shows the rank position of the concept. a) The top 10 concepts with the highest average rank. b) Concepts with the highest rank trend slope. c) Concepts with the lowest rank trend slope. d) The bottom 10 concepts with the lowest average rank.
In level 3, four concepts maintain a stable ranking, with B01.050 (Animals) leading. Concepts E05.481 (In Vitro Techniques), A03.620 (Liver), E01.789 (Prognosis), and N06.850 (Public Health) move among multiple positions, with ‘Prognosis’ tending to climb positions over time. In level 4, only the top three concepts, M01.060.116 (Adult), M01.060.406 (Child), and M01.060.057 (Adolescent), remain stable over time. These concepts continue to generate high research interest. The remaining concepts mainly decrease one position in the ranking.
Figure 4: Top level-3 concept relevance over time for categories (a) C, (b) D, and (c) E.
We computed the rank trend of concepts by adding the difference of two consecutive rank positions of a concept and dividing it by the number of differences, i.e., the number of years minus 1. Thus, we identified the concepts with the highest and lowest slopes. The top 10 upward concepts are shown in Figure 3b, which seem to lose positions over time at levels 2, 3, and 4. The behavior of the rank trend of concepts in these levels is similar. The concepts with the highest slope in level 2 are D06 (Hormones, Hormone Substitutes, and Hormone Antagonists), I01 (Social Sciences), D05 (Macromolecular Substances), and C24 (Occupational Diseases). The concept with the highest slope in level 3 is D01.930 (Thorium Compounds), and in level 4 is D03.383.517 (Oxathiins). The downwards concepts are shown in Figure 3c, which seems to gain positions over time, in contrast to the upward concepts. The concepts with the lowest rank trend slope are G16 (Biological Phenomena) and I02 (Education), both in level 2. In level 3, the concept M01.928 (Vegetarians) has the lowest trend slope, while in level 4, the concept E05.200.312 (COVID-19 Testing) has the lowest slope.
The bottom concepts are shown in Figure 3d. Concepts in level 2 remain stable, while in levels 3, and 4, concepts change positions constantly. On the other hand, bottom concepts in levels 3 and 4 change in a range of hundreds of positions, as in these levels, there are more concepts. Most of the changes occur towards the end of the studied period. The concepts J02.750 (‘Sustenance,’ in level 3) and B03.280.622 (‘Pannus,’ in level 4) seem to gain positions over time. In Figure 3d, right panel, we are only showing the plot since 2017 for visibility sake, as previous years follow the trend of 2017-2018. The MeSH headings of the concepts in Figure 3 can be found in Table 3, Table 4, and Table 5.
We analyzed the top concepts in concept relevance over time in categories C, D, and E at level 2. We mainly found no changes, except in category D, which presented a few changes. Thus, we show in Figure 4 the top concepts in these categories at level 3. In Figure 4a, four of the changing concepts belong to the branches of C12 (Urogenital Diseases) and C14 (Cardiovascular Diseases). The top concepts in Category D at Level 3 exhibit fewer changes than in Level 2. These concepts are D12.776 (Proteins), D12.644 (Peptides), and D12.125 (Amino Acids). In category E, E05.337 (Evaluation Studies as Topic) consistently loses positions over time, and E05.599 (Models, Theoretical) climbs some positions. Concepts E05.481 (In Vitro Techniques) and E01.789 (Prognosis) alternate in the top two positions. On the other hand, E05.581 (Methods) and E05.318 (Epidemiologic Methods) alternate the fourth and fifth positions. The MeSH headings of the concepts in Figure 4 can be found in Table 4.
3.3 Quantitative assessment
In this section, we describe the quantitative assessment of the multiple relevance aspects through the concept evolution task, and retracted papers. We used an ontology evolution dataset for the concept evolution task and identified retracted papers in PubMed. We start this section by describing the datasets, then detail each assessment and present their results.
3.3.1 Datasets for quantitative assessment
Ontology evolution dataset
We created our dataset for concept evolution as previous studies still need to release their datasets. This dataset contains releases where MeSH descriptors are annotated with the type of change, ontological features, and data-driven temporal features. We calculated the latter features from the Data-properties dataset (explained in section 2.1).
A MeSH descriptor (or concept) has evolved if it has a change in the concept’s description (change of description), the addition of new children to the current concept (extension), a change of the concept’s location (move), or the removal of the concept (removal)  [14]. Table 2 shows the distribution of descriptors and changes over releases. We use the ontology versions of MeSH available in BioPortal, which have two versions per year. For example, in 2014, there were the 2014AA and 2014AB releases. When we downloaded the MeSH releases [^5] Downloaded on April 12, 2022, they were available from 2013 to 2021.
Table 2: Distribution of concepts over the MeSH releases from 2014 until 2021. The last two columns show the number of changed concepts in each release.
MeSH
release
Number of
concepts
Did not
evolve
Mean
relevance
Evolved
Mean
relevance
2014AA 27,149 26,833 8.81E-04 316 2.53E-03
2014AB 27,455 27,341 8.73E-04 114 5.40E-03
2015AA 27,455 26,758 8.62E-04 697 2.04E-03
2015AB 27,883 27,879 8.80E-04 4 8.44E-04
2016AA 27,885 26,896 8.40E-04 989 1.96E-03
2016AB 28,472 28,472 8.65E-04 0 -
2017AA 28,472 27,673 8.35E-04 799 1.89E-03
2017AB 28,939 28,939 8.53E-04 0 -
2018AA 28,939 28,318 8.31E-04 621 1.88E-03
2018AB 29,352 29,349 8.43E-04 3 1.91E-03
2019AA 29,351 28,862 8.29E-04 489 1.67E-03
2019AB 29,640 29,640 8.36E-04 0 -
2020AA 29,640 29,279 8.20E-04 361 2.13E-03
2020AB 29,917 29,917 8.30E-04 0 -
2021AA 29,917 29,538 8.08E-04 379 2.51E-03
We used COntoDiff [37] to get the difference between two consecutive MeSH releases and found no changes in 2013. Thus, we excluded this year. We identified which concepts changed in each release and the change a concept has (change of description, extension, move, and removal). We used data until 2021AA, as we used 2021AA and 2021AB to identify the concepts that changed in 2021AA. Evolving concepts represent around 1% of concepts per release, and releases AB have no or very few changes.
Retracted papers in PubMed
We identified the retracted papers in the entire collection of PubMed 2022. Then, we identified which articles in the monthly citation networks (described in section 2.1) were retracted or non-retracted. On average, 800 papers were found to be retracted per month. We then retrieved the MeSH annotations of each article, along with the tree nodes associated with each MeSH descriptor. We obtained a value per metric for each article using these tree nodes by summing their metrics’ values.
Since our citation networks are based on a random sample, one paper may appear in consecutive months. Thus, we obtained an average of each metric per paper per year. We used these values to perform the statistical test.
3.3.2 Assessment in concept evolution
We focused on the concept evolution task to investigate whether the approach proposed in this work could contribute to biomedical terminology evolution. We performed a statistical test of significance using the Mann-Whitney test to determine if there is a significant difference between the relevance of descriptors that changed and descriptors that did not change among releases. We chose this non-parametric test as our data does not follow a normal distribution. For the concept relevance (RRF), we found that when there are evolving descriptors in a release, there is a statistical difference with a p-value less than 0.05, except for release 2015AB.
Figure 5a shows the test results for each relevance aspect in every release. Releases 2016AB, 2017AB, 2019AB, and 2020AB do not contain evolving descriptors. Consequently, more data is needed to compute the statistical test for these releases. Five out of eleven releases exhibit statistically significant differences in usefulness, while in influence and disruptiveness, eight show statistically significant differences. All releases show statistically significant differences in informativeness, but only one release in concept relevance does not demonstrate a statistical difference.
Release Disruptiveness Influence Informativeness Usefulness
Concept relevance
2014AA $p<0.001$ $p<0.001$ $p<0.001$ $p=0.03$ $p<0.001$
2014AB $p=0.25$ $p=0.13$ $p<0.001$ $p<0.001$ $p<0.001$
2015AA $p<0.001$ $p<0.001$ $p<0.001$ $p=0.64$ $p<0.001$
2015AB $p=0.53$ $p=0.56$ $p=0.03$ $p=0.35$ $p=0.29$
2016AA $p<0.001$ $p<0.001$ $p<0.001$ $p=0.005$ $p<0.001$
2016AB - - - - -
2017AA $p<0.001$ $p<0.001$ $p<0.001$ $p=0.11$ $p<0.001$
2017AB - - - - -
2018AA $p<0.001$ $p<0.001$ $p<0.001$ $p=0.22$ $p<0.001$
2018AB $p=0.07$ $p=0.07$ $p=0.04$ $p=0.06$ $p=0.04$
2019AA $p<0.001$ $p<0.001$ $p<0.001$ $p=0.31$ $p<0.001$
2019AB - - - - -
2020AA $p<0.001$ $p<0.001$ $p<0.001$ $p=0.001$ $p<0.001$
2020AB - - - - -
2021AA $p<0.001$ $p<0.001$ $p<0.001$ $p=0.002$ $p<0.001$
Figure 5: Evaluation of the multiple aspects in concept evolution. a) Results of the statistical significance test for each release per aspect using annotations of concept evolution. Green-coded cells correspond to a $p-value<0.05$; otherwise, it is red. b) The top evolving concepts according to each aspect are disruptiveness, influence, informativeness, and usefulness.
Figure 5b shows the concepts that mostly evolve according to each aspect. Some concepts change their position in disruptiveness and influence. Concepts of informativeness and usefulness remain mostly constant. The MeSH headings of concepts in Figure 5b can be found in Table 4.
3.3.3 Assessment in retracted papers
We used the multiple aspects metrics in this study to determine whether there is a significant difference between retracted and non-retracted papers in PubMed. We used the non-parametric Mann-Whitney test since our data does not follow a normal distribution. The results, presented in Figure 6a, revealed a significant difference between retracted and non-retracted papers in concept relevance, using the proposed aspects as the comparison unit, with a p-value less than 0.05. Further, when we performed the test per aspect, we found that there was also a statistical difference in usefulness. In influence, there is a significant difference in five out of eight years, while in disruptiveness and informativeness, seven show statistical differences.
Figure 6b shows the top concepts in retracted papers according to the rank relevance at level 3. Concepts B01.050 (Animals), A11.251 (Cells, Cultured), and E01.789 (Prognosis) lead the rank. Half of the concepts come from branch G, and three belong to branch G04 (Cell Physiological Phenomena). The MeSH headings of the concepts in Figure 6b can be found in Table 4.
Year
Disrupti-
veness
Influence
Informati-
veness
Usefulness
Concept
relevance
2014 $p=0.02$ $p<0.001$ $p<0.001$ $p<0.001$ $p<0.001$
2015 $p=0.68$ $p<0.001$ $p<0.001$ $p<0.001$ $p<0.001$
2016 $p=0.03$ $p=0.002$ $p<0.001$ $p<0.001$ $p<0.001$
2017 $p<0.001$ $p=0.56$ $p<0.001$ $p<0.001$ $p<0.001$
2018 $p<0.001$ $p=0.69$ $p<0.001$ $p<0.001$ $p<0.001$
2019 $p<0.001$ $p=0.38$ $p<0.001$ $p<0.001$ $p<0.001$
2020 $p<0.001$ $p<0.001$ $p=0.15$ $p<0.001$ $p<0.001$
2021 $p<0.001$ $p<0.001$ $p=0.05$ $p<0.001$ $p<0.001$
Figure 6: Evaluation of the multiple aspects in retracted papers. a) Results of the statistical significance test in retracted papers per aspect per year. Green-coded cells correspond to a $p-value<0.05$; otherwise, it is red. b) The top concepts in retracted papers according to the concept relevance at level 3.
4 Discussion
Our study focused on analyzing the evolution of MeSH using a data-driven approach. Leveraging concepts from information theory (i.e., entropy and category utility) and network analysis (i.e., disruptiveness and graph centrality), we investigated different aspects of relevance of MeSH concepts, such as informativeness, usefulness, disruptiveness, and influence. To summarize these four metrics in an overall concept relevance rank, we use the ranking fusion approach - RRF - from information retrieval. In our experiments, we analysed the relevance of MeSH concepts to the PubMed corpus, using the 2022 snapshot annotated with MeSH 2022. Using monthly data from 2014 to 2021, we compute the different relevance aspects, and show how they vary over time. We apply the methodology to two use-cases - concept evolution and retracted articles - and analysed the MeSH concept relevance in these scenarios.
Overall, there is an increasing trend for informativeness and a decreasing trend for usefulness, influence, and disruptiveness. The latter is similar to a recent finding from a study of Park et al., which shows that disruptiveness in life and biomedicine sciences is decreasing  [36]. At the high level of the MeSH categories, our results showed categories C (Diseases)and D (Chemicals and Drugs) are the most relevant for the PubMed collection. While this result is somehow expected from the conceptual level, given the importance of such concepts in life and biomedical sciences, our method is able to quantify it according to different relevance aspects as well as an aggregated relevance rank. In particular, category D ranks first in terms of informativeness and usefulness while category C is the top category regarding influence and disruptiveness. When examining the metrics trends at level 2 of categories C, D, and E (Analytical, Diagnostic and Therapeutic Techniques, and Equipment), they behave similarly to level 1, showing mainly increasing trend informativeness and decreasing trends for the remaining aspects. The correlation analysis of descriptors revealed a weak correlation between informativeness, usefulness, and disruptiveness, demonstrating that indeed they capture different aspects of concept relevance.
When analyzing the top 10 relevant concepts, higher levels tend to be more stable than lower levels due to the propagation of lower-level relevance scores. For example, level 1 appears more stable than level 2. Starting in level 2, i.e., for codes with 3 digits, we notice important changes in the relevance rank over time, with concepts gaining or losing several to many positions during the period analyzed. For example, the concept Chemical Phenomena (G02) falls 4 positions from 2014 to 2021. Levels 3 (6 digits) and 4 (9 digits) show an even more dynamic ranking. In level 3, Prognosis (E01.789) gains 3 positions, while Liver (A03.620) loses 3 positions in the same period. In level 4, most concepts in the top 10 lose one position over time. When examining the top concepts within categories C, D, and E, level 2 remains primarily stable. Among the top concepts within categories C and E at level 3, Cardiovascular Diseases (C14) and Investigative Techniques (E05) are the branches containing more changes, which might reflect changes around the scientific knowledge around those topics. Overall, we notice more concept relevance changes in lower ranks (bottom 10) than in upper ranks (top 10), apart from level 1.
We further evaluated the application of our proposed approach using two analysis use-cases: concept evolution and retracted papers. For the former, we utilized an ontology evolution dataset, containing concept changes from consecutive MeSH versions. Interesting, the Mann-Whitney test confirmed a statistical difference between the relevance rank of MeSH concepts that evolve versus those that remain the same between version. This result suggest that the relevance of a concept could be used to identify candidates for concept update in KOS update. Additionally, our approach successfully discriminates between retracted and non-retracted papers even though the monthly proportion of retracted papers was approximately 0.0002%. In particular, concepts Animals (B01.050), Cells, Cultured (A11.251), Prognosis (E01.789), Age Groups (M01.060), Gene Expression Regulation (G05.308), and Biochemical Phenomena (G02.111) showed sustainable high-relevance to the retracted corpus. While the former 3 concepts also appear in the overall top 10 relevance list, the latter 3 concepts are particular of retracted papers, suggesting potential life and biomedical research areas with issues.
This study has several limitations. First, as we use biomedical data sources, we must have a network of data instances using a KOS, i.e., PubMed and MeSH, to extend this work to other domains. Second, all the metrics were computed using a sample of the actual citation network. This was due to the algorithmic complexity of some of the metrics, which had a prohibitive computational time. Third, metrics are somehow correlated. Within a short period (one release to the other), our method is not sensitive enough to accurately identify all the changes between releases. Fourth, experiments in this work use PubMed 2022, which is annotated with MeSH 2022. More comprehensive experiments are needed to consider more MeSH versions.
5 Conclusions
In this paper, we propose a methodology based on information theory and network analysis to compute the relevance of MeSH concepts for the biomedical corpus indexed by PubMed. By considering multiple relevance aspects of a concept in a data-driven analysis, our approach provides insights into the relevance evolution of MeSH concepts to the PubMed corpus.
In our study, we analyzed the dynamics of MeSH over time and demonstrated the potential of our approach in various tasks, such as concept evolution and identifying retracted papers. We recognize the need for further research to expand our approach to other domains and address the computational complexities associated with some metrics. In the future, we will extend the proposed approach to different domains.
6 CRediT authorship contribution statement
Jenny Copara: Conceptualization, Software, Formal analysis, Investigation, Data Curation, Writing - Original Draft, Visualization. Nona Naderi: Validation, Software, Data Curation, Writing - Review & Editing. Gilles Falquet: Validation, Writing - Review & Editing, Supervision. Douglas Teodoro: Conceptualization, Validation, Formal analysis, Writing - Review & Editing, Supervision, Funding acquisition.
7 Acknowledgment
Part of this work has been funded by the CINECA project (H2020 No 825775) and Innosuisse project funding number 46966.1 IP-ICT.
Appendix A Appendix
Figure 7: Distribution of articles in PubMed from January 1945 to December 2021
Table 3: MeSH headings of concepts in level 2 used in Figures 2 and 3.
Concept MeSH heading Concept MeSH heading
A08 Nervous System D10 Lipids
A11 Cells D12 Amino Acids, Peptides, and Proteins
A13 Animal Structures D26 Pharmaceutical Preparations
A15 Hemic and Immune Systems E01 Diagnosis
A18 Plant Structures E04 Surgical Procedures, Operative
A19 Fungal Structures E05 Investigative Techniques
A20 Bacterial Structures E06 Dentistry
A21 Viral Structures E07 Equipment and Supplies
B02 Archaea G02 Chemical Phenomena
B03 Bacteria G03 Metabolism
B05 Organism Forms G04 Cell Physiological Phenomena
C01 Infections G05 Genetic Phenomena
C04 Neoplasms G09
Circulatory and Respiratory
Physiological Phenomena
C08 Respiratory Tract Diseases G13
Integumentary System Physiological
Phenomena
C09 Otorhinolaryngologic Diseases G15 Plant Physiological Phenomena
C12 Urogenital Diseases G16 Biological Phenomena
C21 Disorders of Environmental Origin I01 Social Sciences
C23
Pathological Conditions, Signs
and Symptoms
I02 Education
C24 Occupational Diseases J02 Food and Beverages
D01 Inorganic Chemicals J03
Non-Medical Public and Private
Facilities
D02 Organic Chemicals N03
Health Care Economics and
Organizations
D05 Macromolecular Substances N04 Health Services Administration
D06
Hormones, Hormone Substitutes,
and Hormone Antagonists
Table 4: MeSH headings of concepts in level 3 are used in Figures 3, 4, 5, and 6.
Concept MeSH heading Concept MeSH heading
A03.620 Liver E02.990 Treatment Switching
A07.521 Glymphatic System E04.625 Prophylactic Surgical Procedures
A11.251 Cells, Cultured E05.196 Chemistry Techniques, Analytical
A20.484 Bacterial Outer Membrane E05.198
Chromophore-Assisted Light
Inactivation
B01.050 Animals E05.200 Clinical Laboratory Techniques
B01.675 Retortamonadidae E05.318 Epidemiologic Methods
B04.352 Fungal Viruses E05.337 Evaluation Studies as Topic
C01.925 Virus Diseases E05.481 In Vitro Techniques
C04.557 Neoplasms by Histologic Type E05.581 Methods
C06.405 Gastrointestinal Diseases E05.599 Models, Theoretical
C10.228 Central Nervous System Diseases E07.123 Capillary Tubing
C12.050
Female Urogenital Diseases
and Pregnancy Complications
F01.145 Behavior
C12.200 Male Urogenital Diseases F01.872 Psychosocial Functioning
C14.280 Heart Diseases F02.830 Psychophysiology
C14.907 Vascular Diseases F04.754 Psychotherapy
C15.378 Hematologic Diseases G02.111 Biochemical Phenomena
C22.232 Disease Models, Animal G02.211 Drug Liberation
C22.557 Mink Viral Enteritis G04.146 Cell Death
C23.888 Signs and Symptoms G04.161 Cell Growth Processes
C24.580 Occupational Stress G04.198 Cell Movement
C26.741 Reinjuries G04.375 Clonal Evolution
D01.268 Elements G04.670 Organelle Size
D01.930 Thorium Compounds G05.158 Clonal Evolution
D01.940 Tungsten Compounds G05.308 Gene Expression Regulation
D02.092 Amines, G06.591 Microbiota
D06.472 Hormones G15.579 Plant Defense Against Herbivory
D08.811 Enzymes I01.880 Sociology
D12.125 Amino Acids I03.743 Stakeholder Participation
D12.644 Peptides I03.833 Temperance Movement
D12.776 Proteins J01.897 Technology
D13.150 Antisense Elements (Genetics) J02.750 Sustenance
D13.695 Nucleotides K01.883 Romanticism
D23.050 Antigens L01.731 Postcards as Topic
D23.383 Immune Checkpoint Proteins M01.060 Age Groups
D27.505 Pharmacologic Actions M01.925 Transplant Recipients
E01.055 Clinical Decision-Making M01.928 Vegetarians
E01.171 Diagnosis, Differential N03.109
Dual MEDICAID MEDICARE
Eligibility
E01.370 Diagnostic Techniques and Procedures N04.590 Patient Care Management
E01.789 Prognosis N06.850 Public Health
Table 5: MeSH headings of concepts in level 4 are used in Figure 3.
Concept MeSH heading Concept MeSH heading
A08.186.211 Brain G03.787.321 Drug Liberation
A11.436.054 A549 Cells G06.591.375 Gastrointestinal Microbiome
B01.650.080 Carnivorous Plant G07.775.813 Salt Stress
B03.280.622 Pannus G08.686.525 Lacteal Elimination
C05.330.663 Metatarsal Valgus G15.568.500 Crassulacean Acid Metabolism
C19.700.419 Hypophysitis G15.610.350 Phytoplasma Disease
D03.383.517 Oxathiins G17.485.500 Deep Learning
D12.776.543 Membrane Proteins I01.738.708 Political Activism
D13.444.154 Cell-Free Nucleic Acids I03.350.062 Compulsive Exercise
D26.310.407
Emtricitabine, Rilpivirine,
Tenofovir Drug Combination
J01.897.564 Printing, Three-Dimensional
D26.310.813
Tobramycin, Dexamethasone
Drug Combination
K01.517.732 Philology
E01.789.800 Treatment Outcome L01.399.375 Data Management
E02.190.204 Auriculotherapy L01.453.233 Data Management
E04.749.500 Robotic Surgical Procedures M01.060.057 Adolescent
E05.200.312 COVID-19 Testing M01.060.116 Adult
E05.318.870 Treatment Switching M01.060.406 Child
F04.754.932 Therapeutic Alliance M01.060.703 Infant
G01.374.661 Kinetics N03.540.089 African Union
G01.910.857 Time Factors N05.300.559
Nurses Improving Care
for Health System Elders
G02.111.490 Kinetics Z01.586.361 Holy Roman Empire